Susan plans to invest $400 at 15% annual interest compounded continuously. When will her investment triple in value?


1
Expert's answer
2022-05-06T03:24:22-0400

If the interest is compounded continuously for t years at a rate of r per year, then the compounded amount is given by:

"A=P\cdot e^{rt},"

P: the principal, amount invested

A: the new balance

t: the time

r: the rate, (in decimal form).


So,

"1200=400\cdot e^{0.15\cdot t},\\
e^{0.15\cdot t}=3,\\
t=\cfrac{\text {ln }3}{0.15}=7.324\text{ years}\approx7\text{ years }4\text{ months.}"